OpenStudy (anonymous):

How do you make this function only accept the domain condition? D: [0,infinity) f(x)=1/(x^2+1)

OpenStudy (anonymous):

hmmm interesting question my fellow studier, if only i had the answer :D

OpenStudy (anonymous):

I'd say you could subtract something "infinitessimally smaller than 1" from 1 in the denominator there.

OpenStudy (anonymous):

true^ bravo lol

OpenStudy (anonymous):

Well I'm given a data table and I have to have a function that works under those conditions. The domain of f(x) is [0,infinity) and the data is x = 1,2,3,4. F(x)= 1/2, 1/5, 1/10, 1/17

OpenStudy (anonymous):

Oh, and you can't take negatives either.

OpenStudy (anonymous):

ughhhhh

ganeshie8 (ganeshie8):

try this : \(f(x) = \frac{1}{(\sqrt{x})^2+1}\)

ganeshie8 (ganeshie8):

its the same function with domain restricted to [0, inf)

OpenStudy (anonymous):

That totally works!